Teo said to Kris, "Give me one peso, and then I'll have as much as you." Kris said to Teo, "No, you give me one peso, and then I'll have twice as much as you." How much did each one have?
Answer by stanbon(75887)

Teo said to Kris, "Give me one peso, and then I'll have as much as you." Kris said to Teo, "No, you give me one peso, and then I'll have twice as much as you." How much did each one have?
-----
Equations:
k-1 = t
t-1 = k/2
-----------------
Rearrange:
k = t+1
k = 2t-2
------
Substitute and solve for "t":
t+1 = 2t-2
t = 3 (Teo has 3 pesos)
----
solve for "k"
k = t + 1
k = 3+1 = 4 (Kris has 4 pesos)
==================================
cheers,
Stan H.